OpenStudy (anonymous):

find the area of the right triangle bounded by the x-axis and y-axis, and the line 3x+5y-15=0.

OpenStudy (ash2326):

 consider the line AB whose equation 3x+5y=15 if y=0 x=5 B(5,0) if x=0 y=3 A(0,3) Area = 1/2* Base* height Area=1/2* OB*AO OB=5 units AO=3 units Area=1/2*5*3=7.5 square unit